OpenStudy (anonymous):

How many liters of oxygen gas at STP are needed to react completely with 8.0 of magnesium?

OpenStudy (anonymous):

2Mg + O2 ---> 2MgO Moles of Mg: 8/ 24 = 1/3 mol Volume of O2 needed: (1/6) * 22.4 = 3.73 L

OpenStudy (anonymous):

3.7 L

OpenStudy (anonymous):

Can you better explain how you got the volume of O2 needed?

OpenStudy (anonymous):

Based the ratio mole that Oxygen = 1/2 of Magnesium!